# Why We Experience Society Differently: Intrinsic Dispositions as Drivers of Ideological Complexity in Adaptive Social Networks

### v2

@@ 4 Conclusions
This study investigated the dynamics of opinion evolution in adaptive social networks by leveraging normalized Lempel-Ziv (nLZ) complexity as a measure of opinion trajectory unpredictability. Through simulations across three experimental settings—pure, dual faction, and heterogeneous—we identified robust, behavior-specific opinion complexity signatures associated with nodes biased toward homophily, neophily, and social conformity. Notably, these signatures were often counterintuitive: homophilic nodes—despite seeking similarity—became increasingly unpredictable; neophilic nodes—despite courting novelty—stabilized early; and conformic nodes followed a U-shaped progression from initial stability to later unpredictability.
A central and novel contribution of this work is the discovery that the unpredictability of an individual’s opinion trajectory is remarkably consistent across varying network compositions, and is instead primarily governed by internal behavioral predispositions. Whether situated in homogeneous or heterogeneous environments, each node type maintained its characteristic nLZ complexity profile. This suggests that intrinsic behavioral factors—such as a preference for similarity, novelty, or conformity—play a dominant role in shaping one’s social experience, often overriding the influence of external social structure.
This insight has broader implications beyond the technical findings. It implies that the complexity of one’s ideological or social trajectory is not merely a reflection of environmental volatility, but rather an emergent property of individual temperament. For example, a homophilic individual may experience growing instability not because of external disruption, but because their behavioral orientation fosters feedback loops that amplify unpredictability. In contrast, a neophilic agent might quickly encounter and assimilate novelty, leading to stabilized opinion dynamics. These patterns reveal that the perceived chaos or cohesion of one’s social experience may be largely self-structured.
From a societal perspective, this finding challenges common assumptions that attribute polarization and ideological volatility primarily to structural or environmental causes. If unpredictability in opinion dynamics arises chiefly from within, then interventions aimed at promoting social stability or cohesion may need to address underlying cognitive and behavioral dispositions, rather than focusing solely on reforming external conditions. This dual approach—targeting both structure and behavior—could yield more effective strategies for mitigating polarization and fragmentation.
Our study also highlights the persistence of individual behavioral “fingerprints” even under strong interaction effects. While heterogeneous environments do induce mild shifts in complexity trends, they do not erase the fundamental patterns tied to each behavioral type. This underscores the resilience and primacy of internal behavioral rules in shaping long-term opinion dynamics.
Finally, the uniqueness of our approach lies not merely in using nLZ complexity, but in focusing on temporal shifts and individualized complexity trajectories in behaviorally diverse networks. Future work could extend this perspective by incorporating complementary complexity measures—such as entropy, fractal dimension, or recurrence quantification—to uncover additional facets of opinion dynamics. More broadly, these results suggest a need for a paradigm shift in opinion dynamics research—one that centers individual behavioral identity as a foundational element in the emergence of social complexity.

### v3

@@ 4 Conclusions
In this study, we examined opinion evolution in an adaptive social network model through the lens of normalized Lempel–Ziv (nLZ) complexity, using it as a measure of the temporal unpredictability of individual opinion trajectories. Across three experimental settings—pure, dual-faction, and heterogeneous networks—we observed behavior-specific complexity patterns associated with nodes biased toward homophily, neophily, and social conformity. These patterns were often counterintuitive within the context of the model: homophilic nodes tended to exhibit increasing unpredictability over time; neophilic nodes displayed stable and bounded complexity despite continual novelty-seeking; and conformic nodes showed a two-phase pattern, with early predictability followed by later growth in unpredictability.
A key empirical result of this work is that the qualitative nLZ complexity patterns associated with different behavioral dispositions persist across diverse network compositions. More importantly, this consistency is complemented by explicit causal intervention tests. By altering the internal behavioral parameters of individual nodes while holding the surrounding social environment fixed, we demonstrate that changes in internal disposition induce systematic and directionally consistent shifts in the symbolic complexity of opinion trajectories. Within the context of this adaptive social network model, these intervention results provide direct evidence that internal behavioral rules exert a causal influence on the unpredictability of individual opinion trajectories, rather than merely correlating with favorable environmental conditions.
Supplementary permutation entropy analyses for Scenarios 1 and 3 confirm the presence of non-random ordinal structure but do not distinguish behavioral factions, reinforcing that the observed differences arise from long-range temporal organization captured by nLZ complexity (Figures S8-S11).
Within the limited but informative scope of this model, the results suggest an intriguing perspective on individual-level social experience. The complexity or volatility of an agent’s opinion trajectory appears to be closely linked to its own interaction rules, rather than being determined solely by the composition of the surrounding network. In this sense, the model raises the possibility that perceived instability or predictability in social experience may emerge endogenously from behavioral tendencies, rather than exclusively from external social conditions. While caution is essential when extrapolating from stylized models to real social systems, this perspective resonates with broader themes in psychology and social science concerning the interplay between individual dispositions and environmental influence. Rather than framing social instability purely as a consequence of external structure, the results highlight the potential importance of internal behavioral orientations in shaping subjective experience. The adaptive social network model studied here does not claim to capture human behavior in full complexity, but it offers a tractable setting in which such ideas can be systematically explored and refined.
Finally, the contribution of this work lies not only in the use of nLZ complexity, but in its emphasis on temporal, individual-level dynamical signatures within behaviorally heterogeneous adaptive networks. By combining observational analyses with targeted internal interventions, we demonstrate how behavioral dispositions can be causally linked to differences in opinion trajectory complexity within this modeling framework. Future work could extend this approach by incorporating complementary dynamical and information-theoretic measures—such as entropy-based metrics, recurrence quantification analysis, or fractal characteristics—to provide additional lenses on opinion dynamics, as each tool captures distinct aspects of temporal structure. At the same time, the explanations offered for the observed complexity trends remain phenomenological and conjectural; while causality is established at the level of internal parameters, a rigorous mechanistic derivation linking nLZ patterns to specific adaptive network processes—such as link weight evolution, clustering dynamics, or rewiring activity—remains an important direction for future work. More broadly, these results point toward a research direction in opinion dynamics that treats individual behavioral identity as a meaningful contributor to social complexity, while remaining attentive to the assumptions and limitations inherent in simplified models of social systems.